# Towards a more realistic evaluation of machine learning models for bearing fault diagnosis

### v1

@@ 1 Introduction
Despite the significant advancements offered by ML, its application requires a careful methodology to ensure models generalize reliably to real-world scenarios. A critical methodological pitfall in ML-based science is data leakage [2]. Data leakage is defined as a spurious relationship between independent variables and the target variable, arising from flaws in data collection, sampling, or preprocessing [3]. Such an artifact, not present in the true data distribution, typically leads to overoptimistic estimates of model performance. This phenomenon is a major source of error in ML applications, often causing published models to fail when deployed in practical settings.
The problem of data leakage is not unique to bearing fault diagnosis; it is a widespread failure mode in ML-based science, impacting at least 294 papers across 17 scientific fields [3]. For instance, in medicine, improper handling of patient data can lead to leakage if samples from the same patient are used in both training and test sets. Cases have been reported where models included features that were effectively proxies for the outcome, such as the use of anti-hypertensive drugs to predict hypertension or antibiotics to predict sepsis, leading to artificially inflated performance.
@@
Our observations indicate that the problem of data leakage is prevalent in the area of bearing fault diagnosis. Many studies have failed to properly partition their datasets, leading to information leakage and consequently, over-optimistic performance estimates that do not hold in real-world scenarios. For example, studies assigning waveform recordings from the same bearing to both training and test partitions have reported inflated performance. Early work by [4] highlighted this “similarity bias” in ML research using vibration data, finding that nearly all reviewed studies, including 40 out of 41 using the widely adopted Case Western Reserve University (CWRU) dataset, employed experimental designs susceptible to this bias. Recent publications continue to show overestimated results due to similarity bias. Comprehensive reviews in vibration-based intelligent fault diagnosis have also largely overlooked this crucial issue.
Previous efforts have attempted to mitigate data leakage in bearing fault diagnosis. Hendriks et al. [5], for instance, proposed a dataset split for CWRU based on fault size rather than motor load, which significantly reduced the inflated classification accuracy observed in earlier studies. Similarly, [6] explored alternative data splits, demonstrating that assigning vibration data from a given bearing across training and evaluation splits leads to over-optimistic performance estimates. Their work showed notable performance drops (e.g., Naive Bayes accuracy decreased from 85.8% to 69.5% when moving from a random split with leakage to a bearing-wise split) when leakage was addressed. Despite these valuable contributions, a common limitation remained on the CWRU dataset: their multi-class formulations can not entirely eliminate leakage from the healthy class, as this dataset contains only one healthy bearing configuration. Additionally, since the CWRU dataset contains only faulty bearings with a single fault, their multi-class approach cannot account for scenarios where multiple fault types might occur simultaneously.
Given these challenges, this paper aims to further advance the reliable development and deployment of ML models for bearing fault diagnosis. We propose a novel methodology that rigorously addresses data leakage and class imbalance, particularly for datasets with limited healthy bearing data. Our approach formulates the problem as a binary multi-label classification for each sensor location (e.g., drive end and fan end), enabling the detection of the presence or absence of each fault type (e.g., inner, outer, ball). This formulation enables a more precise evaluation by using prevalence-independent metrics such as the Receiver Operating Characteristic (ROC) curve and the Area Under the Receiver Operating Characteristic curve (AUROC). Furthermore, it provides a more realistic representation of real-world conditions where multiple fault types could coexist. Our contributions include:
  • Demonstrating the significant impact of data leakage on classification accuracy across different train/test splits, revealing that models that appear to perform well with leakage may exhibit a substantial drop in performance (e.g., over 40% accuracy drop in some cases) when evaluated without it.
@@
  • Applying the proposed methodology on widely-used vibration datasets, such as CWRU, Paderborn University (PU), and University of Ottawa (UORED-VAFCLS) datasets. Our results reveal the significant impact of bearing diversity on model generalization and demonstrate that the optimal choice between deep and shallow learning models is highly dataset-dependent.
@@ 2.1 Basic concepts on supervised learning
$R(f_{\theta})=\mathbb{E}_{(x,y)\sim P(X,Y)}[\mathcal{L}(f_{\theta}(x),y)].$
@@
$R_{emp}(f_{\theta})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i}).$
@@
$\theta^{*}=\arg\min_{\theta}R_{emp}(f_{\theta}).$
@@
$\lambda^{*}=\arg\min_{\lambda}\left(\frac{1}{|D_{val}|}\sum_{(x,y)\in D_{val}}\mathcal{L}(f_{\theta^{*}(\lambda)}(x),y)\right).$
@@ 3.1.1 Data Leakage in Bearing Fault Diagnosis
A condition-wise split happens when signals with different machine conditions are divided between training and testing. A condition is usually represented by the combination of machine configurations, such as the load and rotation speed, for instance. It can also be represented by a severity level in a given bearing, as in the UORED-VAFCLS dataset, where two fault severity levels are available for every bearing. This type of data splitting is commonly used in bearing fault diagnosis, as it is generally understood that a realistic evaluation methodology requires diverse machine configurations to be separated between training and testing. While some may argue that such variation justifies this split, it is conceivable that some intrinsic signature of the bearing remains present in both sets, allowing the model to exploit identity-specific features rather than learning robust patterns.
@@
An even more severe form of data leakage arises from repetition-wise splits (also referred to as run-to-run in [13]), where signals acquired under the same configuration are divided between training and testing. Since signals captured under identical setups typically share similar characteristics, this setup enables models to memorize the signals rather than learn to diagnose faults accurately. In Section 6.2, we show that in the PU dataset, all splits with bearing-level leakage produce overoptimistic results, with repetition- and segment-wise splits reaching almost 100% accuracy.
@@
All these cases result in evaluations that are fundamentally closer to measuring memorization (e.g. the training performance) rather than generalization. To preclude such pitfalls, we advocate for a strict bearing-wise split, in which the physical bearings used for training and testing are mutually exclusive. This approach ensures the model is assessed on its ability to generalize to unseen components, preventing it from leveraging identity-specific artifacts [^2] If samples from the same bearing but with different labels (such as when a healthy bearing is subject to accelerated lifetime tests and develops a fault) are separated into training and test sets, this would not characterize bearing-level leakage, since the spurious correlation between them would not favor the model’s decision towards the correct class. Instead, it would constitute a more challenging evaluation process, which tests whether the model has learned the correct feature representation rather than relying on bearing signatures.. Ultimately, this promotes the development of models that identify universal fault characteristics, a goal best supported by datasets that include multiple physical bearings per fault category. It is important to note that even with bearing-wise splits, leakage may still occur due to spurious correlations learned from poor data collection practices. For example, if all fault signals in a dataset contain additional interference absent from healthy signals, the model might rely on that interference for classification, thus enabling memorization.
@@ 3.2 A Toy Example
First, we established a theoretical performance ceiling for the classifier, as shown in A. In summary, computing the mean of the 3 fault predictive features and applying a threshold of $a_{f}/2$ gives the maximum achievable accuracy of $90.30\%$. We then evaluated models with two distinct test set configurations: the first consisted of samples generated from the same bearings used in the training set, while the second comprised samples generated from different bearings.
Using Logistic Regression and Decision Tree classifiers to assess the impact of model capacity, our results (Figure 1) show that test sets containing samples from bearings included in the training set yield overly optimistic performance metrics for both classifiers, even exceeding the theoretical maximum classification accuracy. Due to data leakage, the results misleadingly suggest that Logistic Regression is the superior model; however, under the valid test set, the Decision Tree actually demonstrates better performance when 4 or more training bearings are used. This highlights that results influenced by data leakage are unreliable for informed decision-making. Additionally, increasing the number of bearings in training improves performance on the leakage-free test set, while it also mitigates the impact of leakage on both models, suggesting it learns to prioritize the correct predictive features. These findings underscore the critical importance of a rigorous validation methodology to prevent misinterpretation of a model’s true generalization capabilities.

### v2

@@ 1 Introduction
Despite the significant advancements offered by ML, its application requires a careful methodology to ensure models generalize reliably to real-world scenarios. A critical methodological pitfall in ML-based science is data leakage [2]. Data leakage is defined as a spurious relationship between independent variables and the target variable, arising from flaws in data collection, sampling, or preprocessing [3]. Such an artifact, not present in the true data distribution, typically leads to overoptimistic estimates of model performance. This phenomenon is a major source of error in ML applications, often causing published models to fail when deployed in practical settings, impacting at least 294 papers across 17 scientific fields [3]. For instance, in medicine, improper handling of patient data can lead to leakage if samples from the same patient are used in both training and test sets. Cases have been reported where models included features that were effectively proxies for the outcome, such as the use of anti-hypertensive drugs to predict hypertension or antibiotics to predict sepsis, leading to artificially inflated performance.
@@
Our observations indicate that data leakage remains a prevalent issue in the field of bearing fault diagnosis. Numerous studies fail to partition datasets correctly, resulting in information leakage and, consequently, over-optimistic performance estimates that do not hold in real-world scenarios. For instance, studies that assign waveform recordings from the same bearing to both training and test partitions have consistently reported inflated performance. Early work by [4, 5] highlighted this “similarity bias” in machine learning research utilizing vibration data, revealing that nearly all reviewed studies (published between 2008 and 2020), including 40 out of 41 using the widely adopted Case Western Reserve University (CWRU) dataset, employed experimental designs susceptible to this bias. These findings are corroborated by [6], who examined 55 papers published between 2020 and 2024 in the bearing fault diagnosis domain, finding that only six employed rigorous data-splitting methodologies. In addition to these previous studies, the present work contributes an investigation of 18 papers published in 2025, identifying that data leakage persists in the majority of works in the field, leading to overestimated results (Section 3.3).
To achieve a more realistic evaluation, this paper advocates for bearing-wise splitting, which consists of ensuring that all data originating from the same bearing is assigned exclusively to a single (train or test) partition. This strategy is necessary to prevent data leakage, since otherwise a model may learn bearing-specific artifacts that are spuriously correlated with the target variable, leading to overly optimistic results. In other words, the model may simply memorize the bearing identity and its associated fault label, rather than learning robust fault signatures that generalize to unseen bearings.
Establishing that conventional (non-bearing-wise) splitting strategies indeed lead to unrealistic results—and thus should be avoided in future work—requires one to show that using a bearing-wise split causes a significant performance gap in an otherwise identical setup. This path was taken by [6, 7, 8, 9], which have observed accuracy dropping from near 100% to around 40%-60% depending on the specific setup. However, [7, 8] proposed splitting data by fault size, which does not necessarily correspond to a bearing-wise split (see Section 3.4), while [7, 9] could not entirely eliminate leakage with respect to the healthy class in the CWRU dataset, as this dataset contains a single healthy bearing configuration.
Additionally, none of these works controlled an important confounding factor in their experimental design: the number of training bearings. Specifically, when naively changing from a traditional split (where all available bearings are used for training) to a bearing-wise split (where the bearings used for testing are not included in the training set), the number of bearings seen during training is reduced. As is well-known in machine learning literature and corroborated by our experiments with bearing data, the diversity of training data (not just the raw number of samples) is an important driver of model performance. Thus, it is conceivable that the aforementioned performance gap could arise simply due to a reduced training diversity. To convincingly show that this is not the case, in this paper we perform controlled experiments where the exact trained model is kept fixed and only the test dataset is changed based on the splitting strategy. To the best of our knowledge, this is the first paper to present such experiments, through which we hope to convince readers that using an appropriate splitting procedure is strictly necessary to produce valid results.
An alternative approach to completely eliminating data leakage is conducting inter-testbench experiments, typically framed within a cross-domain or domain generalization context. In this formulation, the testbenches used for training and testing are completely isolated; however, this introduces the significant challenge of domain generalization. We suspect that the bearing fault diagnosis literature has largely treated intra-domain classification as a solved problem, prompting a shift toward cross-domain research. Research by [10] indicates that many studies define a domain as a combination of operating conditions (e.g., load, rotation speed, torque) rather than distinct physical bearings. This perspective has led to various data splitting methodologies within single datasets, where data is partitioned according to these conditions. A limited number of studies have addressed the inter-testbench scenario by isolating datasets, such as [11], [12], and [13]. Although these approaches effectively eliminate data leakage, domain generalization remains a complex challenge, as target datasets often exhibit disparate feature distributions, necessitating specialized techniques to extract domain-invariant features. The present work, however, focuses on a simpler yet fundamentally critical problem: training and testing within a single test bench.
This paper aims to further advance the reliable development and deployment of ML models for bearing fault diagnosis. We propose a novel methodology that rigorously addresses data leakage and class imbalance in a intra-dataset scenario, particularly for datasets with limited healthy bearing data. Our approach formulates the problem as a binary multi-label classification for each sensor location (e.g., drive end and fan end), enabling the detection of the presence or absence of each fault type (e.g., inner, outer, ball). This formulation specifically addresses the disadvantages of multiclass accuracy, which serves as a poor proxy for real-world performance. Because accuracy treats all misclassifications as equal, it fails to distinguish between “false alarm” (False Positive) and a “missed detection” (False Negative), the latter being significantly more costly in industrial maintenance. Furthermore, in the presence of class imbalance, a model can achieve high accuracy by simply predicting the majority class, effectively “hiding” its inability to identify rare but critical fault states. Another issue caused by the multiclass formulation is the inability of detecting co-occurring faults. While it is possible to create classes for combined faults, it is often unpractical and most public datasets contain few examples of those cases. By treating faults as independent binary problems, we can accommodate co-occurring defects and utilize faulty signals from one label as true negatives for others—a significant advantage when healthy data is scarce, such as in the CWRU dataset. In this formulation, we can also utilize prevalence-independent metrics such as the Area Under the Receiver Operating Characteristic curve (AUROC). This allows for a more realistic representation of real-world conditions where decision thresholds must be precisely tuned to prioritize sensitivity or specificity based on specific safety and operational requirements.
In summary, our contributions include:
  • Designing experiments that isolate data leakage from other confounding factors using both synthetic and real data. By fixing the trained model and comparing a test set containing signals from bearings seen during training (leaked) against another test set composed entirely of unseen bearings, we show that performance degradation is directly attributable to data leakage rather than reduced training data diversity.
@@
  • Applying the proposed methodology on widely-used vibration datasets, such as CWRU, Paderborn University (PU) and University of Ottawa (UORED-VAFCLS) datasets. Our results reveal the significant impact of bearing diversity on model generalization and demonstrate that the optimal choice between deep and shallow learning models is highly dataset-dependent.
@@ 2.1 Basic concepts on supervised learning
$R(f_{\theta})=\mathbb{E}_{(x,y)\sim P(X,Y)}[\mathcal{L}(f_{\theta}(x),y)].$ (1)
@@
$R_{emp}(f_{\theta})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{\theta}(x_{i}),y_{i}).$ (2)
@@
$\theta^{*}=\arg\min_{\theta}R_{emp}(f_{\theta}).$ (3)
@@
$\lambda^{*}=\arg\min_{\lambda}\left(\frac{1}{|D_{val}|}\sum_{(x,y)\in D_{val}}\mathcal{L}(f_{\theta^{*}(\lambda)}(x),y)\right).$ (4)
@@
2.2 Evaluation metrics
Consider now a classification model $f:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{Y}=\{1,\ldots,K\}$ is a finite set and $K$ is the number of classes. Let $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ be random variables whose joint distribution is given by $P(X,Y)$. The primary objective of a classification model is to correctly map inputs to their respective categories, an ability that can be evaluated using quantitative performance metrics. The most widely used metric, accuracy, measures the proportion of correctly classified instances and is defined as
$\text{Acc}(f)=P[f(X)=Y].$ (5)
However, this metric is fundamentally limited by its sensitivity to class prevalence [16]. In problems with significant class imbalance, accuracy becomes misleading, as a model can achieve a deceptively high score by simply predicting the majority class. To address this, metrics more robust to class imbalance are required, such as the balanced accuracy, defined as the arithmetic mean of the per-class recall:
$\text{Bacc}(f)=\frac{1}{K}\sum_{k=1}^{K}P[f(X)=k)|Y=k].$ (6)
For binary classification, where $\mathcal{Y}=\{0,1\}$ and $K=2$, the balanced accuracy can be expressed as
$\text{Bacc}(f)=\frac{\text{TNR}+\text{TPR}}{2}=1-\frac{\text{FPR}+\text{FNR}}{2}$ (7)
where
TNR $=P[f(X)=0)|Y=0]$ (8)
TPR $=P[f(X)=1)|Y=1]$ (9)
FPR $=P[f(X)=1)|Y=0]=1-\text{TNR}$ (10)
FNR $=P[f(X)=0)|Y=1]=1-\text{TPR}$ (11)
which stand for True Negative Rate, True Positive Rate, False Positive Rate and False Negative Rate, respectively. Note that, in this case, accuracy can be equivalently expressed as
$\text{Acc}(f)=(1-p)\cdot\text{TNR}+p\cdot\text{TPR}=1-(1-p)\cdot\text{FPR}-p\cdot\text{FNR},\qquad p=P[Y=1]$ (12)
highlighting its dependence on the prevalence $p$. For instance, a classifier $f(x)=0$ that always predicts the negative class achieves accuracy 1p1-p (while $\text{FPR}=0$ and $\text{FNR}=1$). Ideally, we would like both error metrics FPR and FNR to be close to zero.
Often, in binary classification, we have access to a family of classifiers $f_{\tau}$ parameterized by a threshold $\tau$, specifically, $f_{\tau}(x)=\mathbf{1}[s(x)\geq\tau]$, where $s:\mathcal{X}\to\mathbb{R}$ is a scoring function and $\mathbf{1}(\cdot)$ denotes the indicator function. In this case, we are not limited to a specific pair (FPR, FNR) of error metrics; instead, by sweeping over $\tau$, we can arbitrarily choose a desired operating point. The tradeoff between the achievable (FPR, FNR)—or, more commonly, between (FPR, TPR)—for all possible $\tau$ is known as the Receiver Operating Characteristic (ROC) curve [17].
As the ROC curve represents classifier behavior across a continuum of decision thresholds, it is often convenient to summarize this information into a single scalar value. The Area Under the ROC Curve (AUROC) provides such a summary by computing the integral of the TPR $\times$ FPR curve over the entire $[0,1]$ range. In particular, the AUROC of a perfect classifier equals 1, while that of a random classifier equals 0.5. Because it is computed over all operating points, the AUROC does not depend on the selection of a specific decision threshold. Furthermore, as it is derived from conditional rates rather than absolute class frequencies, the AUROC is invariant to class prevalence.
It is straightforward to extend these definitions to multi-label classification. In binary multi-label classification with $L$ labels, one deals with $L$ independent binary classifiers. Each classifier yields its own score function and corresponding ROC curve, resulting in a set of label-specific AUROC values. To obtain a single representative measure of performance across all $L$ labels during model development, these AUROC values can be aggregated via macro-averaging. The resulting Macro AUROC is defined as the arithmetic mean of the individual AUROC scores:
$\text{Macro AUROC}=\frac{1}{L}\cdot\sum_{i=1}^{L}\text{AUROC}_{i}.$ (13)
This aggregation assigns equal weight to each label and provides a concise summary of overall model behavior, while the individual ROC curves retain their relevance for label-specific analysis and threshold selection.
@@ 3.1.1 Data Leakage in Bearing Fault Diagnosis
A condition-wise split happens when signals with different machine conditions are divided between training and testing. A condition is usually represented by the combination of machine configurations, such as the load and rotation speed, for instance. Alternatively, a condition may be defined in terms of the fault severity level of a given bearing, as in the UORED-VAFCLS dataset, where two distinct severity levels are provided for each bearing. This type of data splitting is commonly used in bearing fault diagnosis, as it is generally understood that a realistic evaluation methodology requires diverse machine configurations to be separated between training and testing. While some may argue that such variation justifies this split, it is conceivable that some intrinsic signature of the bearing remains present in both sets, allowing the model to exploit identity-specific features rather than learning robust patterns.
@@
An even more severe form of data leakage arises from repetition-wise splits (also referred to as run-to-run in [6]), where signals acquired under the same configuration are divided between training and testing. Since signals captured under identical setups typically share similar characteristics, this setup encourages models to memorize signal-specific patterns rather than learn features that generalize to fault diagnosis. In Section 6.2, we show that in the PU dataset, all splits with bearing-level leakage produce overoptimistic results, with repetition- and segment-wise splits reaching almost 100% accuracy.
@@
All these cases result in evaluations that are fundamentally closer to measuring memorization (e.g. the training performance) rather than generalization. To preclude such pitfalls, we advocate for a strict bearing-wise split, in which the physical bearings used for training and testing are mutually exclusive. This approach ensures the model is assessed on its ability to generalize to unseen components, preventing it from leveraging identity-specific artifacts [^2] If samples from the same bearing but assigned different labels (for example, when a healthy bearing is subjected to accelerated lifetime testing and later develops a fault) are split between training and test sets, this configuration does not characterize bearing-level leakage, because the resulting correlations do not inherently bias the model toward the correct class. Rather, it constitutes a more challenging evaluation scenario, as it assesses whether the model has learned meaningful feature representations rather than relying on bearing-specific signatures.. Ultimately, this promotes the development of models that identify universal fault characteristics, a goal best supported by datasets that include multiple physical bearings per fault category. It is important to note that even with bearing-wise splits, leakage may still occur due to spurious correlations learned from poor data collection practices. For example, if all fault signals in a dataset contain additional interference absent from healthy signals, the model might rely on that interference for classification, thus enabling memorization.
@@ 3.2 A Toy Example
First, we established a theoretical performance ceiling for the classifier, as shown in Appendix A. In summary, computing the mean of the 3 fault predictive features and applying a threshold of $a_{f}/2$ gives the maximum achievable accuracy of $90.30\%$. We then evaluated models with two distinct test set configurations: the first consisted of samples generated from the same bearings used in the training set, while the second comprised samples generated from different bearings.
Using Logistic Regression and Decision Tree classifiers to assess the impact of model capacity, our results in Figure 1 show that test sets containing samples from bearings included in the training set yield overly optimistic performance metrics for both classifiers, even exceeding the theoretical maximum classification accuracy. Due to data leakage, the results misleadingly suggest that Logistic Regression is the superior model; however, under the valid test set, the Decision Tree actually demonstrates better performance when 4 or more training bearings are used. This highlights that results influenced by data leakage are unreliable for informed decision-making. Additionally, increasing the number of bearings in training improves performance on the leakage-free test set, while it also mitigates the impact of leakage on both models, suggesting it learns to prioritize the correct predictive features. These findings underscore the critical importance of a rigorous validation methodology to prevent misinterpretation of a model’s true generalization capabilities.
@@
3.3 Prevalence of data leakage in bearing fault diagnosis
As mentioned in the introduction, data leakage is widespread in the literature on bearing fault diagnosis, with previous works [4, 6] reporting data splitting issues in over 90% of published papers. To complement and update these findings, we conducted an investigation of papers published in 2025 using a similar methodology. First, we searched the Mechanical Systems and Signal Processing journal database using the queries “bearing fault diagnosis” and “machine learning”, which yielded 195 published papers from January 2025 up to December 2025. We then randomly selected 10 papers that focus on the problem of training and testing within a single dataset under a supervised machine learning context. Second, we extended our search to include papers from other journals within the same scope, compiling an additional 8 papers.
Table 1 lists these 18 papers, detailing whether they specified the train-test partitioning and, if so, how it was performed. We found that 8 of the 18 papers used a random splitting strategy, while 9 detailed a condition-wise split. The remaining paper did not mention any partitioning methodology. The most common condition considered for partitioning was the load, although other papers considered rotation speed or level of noise. Our conclusion is that despite an increase in the number of articles that detail their train-test partitioning methodology, data leakage remains a prevalent issue.
Table 1: Sample of papers published in 2025 that propose or apply machine learning techniques on bearing fault diagnosis datasets.
Paper Split type Journal/Conference Datasets Used Results
[21] Random Other Other >98.8%
[22] Random MSSP Other >99.75%
[23] Random MSSP CWRU, MFPT + Others 100%
[24] Condition-wise MSSP Others 100%
[25] Condition-wise MSSP UORED-VAFCLS, XJTU-SQV 100%
[26] Condition-wise MSSP CWRU, MFPT, PU 100%
[27] Condition-wise MSSP UORED-VAFCLS, HIT 100%
[28] Condition-wise MSSP PU, JNU, HIT >99.2%
[29] Condition-wise Other CWRU >99.3%
[30] Random Other CWRU >99.9%
[31] Random Other CWRU, XJTU-SY >99.5%
[32] Condition-wise Other CWRU >98.5%
[33] Not detailed MSSP CWRU + Other >94%
[34] Random Other Other >99.3%
[35] Condition-wise Other CWRU + Other >99.7%
[36]
Condition-wise
MSSP UORED-VAFCLS + Other >99.9%
[37] Random MSSP Other >98%
[38] Random Other CWRU, PU >97%